# Parameter-related strong convergence rate and polynomial stability of a Euler’s type method for time-changed stochastic differential equations

### v1

@@ Abstract
A Euler’s type method with the equidistant step size is proposed for a class of time-changed stochastic differential equations driven by the multiplicative noise and the strong convergence rate that is related to the parameter of the time changing process is obtained. Such a observation of the convergence rate is significantly different from those existing results that employ methods with the random step size. The polynomial stability in the mean square sense of the numerical method is also studied, which is in line with the asymptotic behavior of the underlying equation. Numerical simulations are provided to demonstrate the theoretical results.
@@
time-changed stochastic differential equations, equidistant-step Euler-Maruyama method, strong convergence, stability
@@ 1 Introduction
where $\mathcal{L}_{\text{FP}}=-\frac{\partial}{\partial x}\left(f(x)\cdot\right)+\frac{1}{2}\frac{\partial^{2}}{\partial x^{2}}\left(g(x)^{2}\cdot\right)$. In contrast, when considering subdiffusion phenomena via time-changed SDEs, as rigorously established by [5, 4], the probability density function of a time-changed process $X(E(t))$ (to be formally de?ned subsequently) is governed by a fractional Fokker-Planck equation(FFPE)
@@
Given the analytical complexity of time-changed SDEs, numerical methods play a crucial role in their practical implementation. The first systematic investigation of numerical methods for time-changed SDEs was conducted by Jum and Kobayashi [6], who considered the following time-changed SDEs
@@
Following this pioneering work, Jin and Kobayashi [7] investigated a class of time-changed SDEs where the inverse subordinator $E(t)$ in the drift and diffusion coefficients of eq. 1 is replaced with ordinary time $t$. In this modified framework, the duality principle no longer applies, necessitating the development of alternative analytical techniques. Using a Gronwall-like inequality specific to time-changed integrals, they demonstrated that the EM method maintains a strong convergence rate of $1/2$ in this setting as well.
@@
Stability of stochastic systems is also an active research area[14]. Most existing studies of stability for time-changed SDEs focused on the following form
$\mathrm{d}X(t)=h(t,E(t),X(t))\mathrm{d}t+f(t,E(t),X(t))\mathrm{d}E(t)+g(t,E(t),X(t))\mathrm{d}B(E(t)).$ (3)
Wu [15] established stochastic stability, stochastic asymptotic stability, and global stochastic asymptotic stability of time-changed SDEs. Zhu et al. [16] demonstrated almost sure exponential stability under appropriate coefficient conditions, while subsequent work by Zhu et al. [17] established exponential stability, revealing the term $h(t,E(t),X(t))\mathrm{d}t$ in eq. 3 determining these stability properties. Furthermore, the stability theory for time-changed stochastic systems has been extended to more general noise structures by Nane and Ni [18, 19], who established comprehensive stability results for SDEs driven by time-changed Lévy noises. This direction was further developed by Yin et al. [20], who investigated the stability properties of such systems under impulsive effects.
However, as noted in Remark 2.13 of [19], the inverse subordinator $E(t)$, which will be formally introduced in Section 2, exhibits asymptotically slower growth compared to ordinary time $t$. Specifically, it satisfies the asymptotic relation $\lim_{t\to\infty}\frac{E(t)}{t}=0$ almost surely. Therefore, the term $h(t,E(t),X(t))\mathrm{d}t$ tends to dominate the system dynamics, potentially masking the distinctive effects introduced by the time-change mechanism. To isolate and analyze more precisely the impact of time-change on stability properties, Deng and Liu [8] and Liu [21] studied the following time-changed SDEs
$\mathrm{d}X(t)=f(t,E(t),X(t))\mathrm{d}E(t)+g(t,E(t),X(t))\mathrm{d}B(E(t)),$ (4)
in which the system dynamics are governed exclusively by the time-changed components.
Within this more focused framework, Deng and Liu [8] established a profound connection between the polynomial stability index of time-changed SDEs and the stability parameter of the subordinator. Their analysis extended to numerical methods as well, demonstrating polynomial stability for the backward EM scheme. Complementing these findings, Liu [21] investigated polynomial stability under polynomial growth conditions, confirming that this stability property is also fundamentally characterized by the stability parameter of the subordinator.
For analytical simplicity in this paper, we concentrate on a simplified form of (4), specifically
@@
where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times m}$ are coefficient functions satisfying global Lipschitz and linear growth conditions, $E(t)$ denotes the inverse subordinator, and $B(E(t))$ represents the $m$-dimensional time-changed standard Brownian motion.
Our contributions are threefold
@@
  3. 3.
    We conduct a comprehensive stability analysis of the numerical scheme, deriving sufficient conditions for asymptotic mean-square polynomial stability with decay rate proportional to $t^{-\alpha}$, thus establishing a direct link between the memory effects induced by the time-change process and the long-term behavior of the numerical solution.
The remainder of this paper is organized as follows Section 2 introduces the mathematical preliminaries, including key properties of subordinators and their inverses, and establishes important lemmas used in our subsequent analysis. Section 3 presents the fundamental assumptions, develops the equidistant-step EM scheme, and proves our main convergence theorem. Section 4 investigates the stability properties of both the exact and numerical solutions. Section 5 presents numerical experiments that validate our theoretical findings and illustrate the effectiveness of the proposed method.
@@ 2 Preliminaries
To numerically simulate the inverse subordinator $E$, we employ the method developed by Magdziarz [22, 23]. Given a time horizon $T>0$ and an equidistant step size $\delta>0$, we first generate a sample path of $D$ over $[0,T]$ by setting $D(0)=0$ and recursively computing
@@
The equidistant-step EM method presented in eq. 6 has been previously utilized by Carnaffan and Kawai [24] in their investigation of anomalous diffusion processes, which provide stochastic representations for multidimensional FFPE. Their approach employed sophisticated techniques from Malliavin calculus to derive unbiased density formulas for the probability density functions of these anomalous diffusion processes. This methodology yielded remarkably accurate numerical approximations that successfully captured essential characteristics of the underlying processes, including the distinctive cusp formations observed in density profiles. Despite these promising applications, a rigorous mathematical analysis concerning the convergence properties and stability characteristics of this numerical scheme in the context of time-changed SDEs has not been thoroughly established in the literature. The present work endeavors to fill this significant gap in the theoretical understanding of numerical methods.
@@
The following lemma, which can also be found in [25], provides precise bounds on the expected values of multiple integrals involving the inverse subordinator. This result constitutes the cornerstone of our convergence rate analysis, establishing the direct relationship between the time step size and the expectation of multiple integrals driven by increments of $E(t)$.
@@
Let $E$ be the inverse process of a subordinator $D$ with stability index $\alpha\in(0,1)$. For any fixed $\varepsilon>0$ and $a,b$ satisfying $\varepsilon<a<b$, and for any positive integer $n$, the following bound holds
$\mathbb{E}\left[\left(\int_{a}^{b}\mathrm{d}E(t)\right)^{n}\right]\leq C(b-a)^{1+(n-1)\alpha},$
where $C>0$ is a constant independent of bab-a.
Proof.
The proof follows from similar arguments as in [25]. For completeness, we provide a detailed proof in A. ∎
While Lemma 3 provides a general bound for arbitrary powers $n$, it is instructive to examine the special case $n=1$ more closely. Below, we derive an explicit formula for the expected increment of the inverse $\alpha$-stable subordinator over an arbitrary interval $[a,b]$ with $0<a<b$, which complements and illuminates the general result.
Utilizing the well-established moment formula $\mathbb{E}\left[E^{n}(t)\right]=\frac{n!t^{n\alpha}}{\Gamma(1+n\alpha)}$ for $n\geq 1$, we derive the first moment of the increment as follows
$\mathbb{E}\left[E(b)-E(a)\right]$ $=\mathbb{E}\left[E(b)\right]-\mathbb{E}\left[E(a)\right]=\frac{b^{\alpha}-a^{\alpha}}{\Gamma(1+\alpha)}.$
By applying the mean value theorem for differentiable functions, we can identify a point $\xi\in(a,b)$ such that
$b^{\alpha}-a^{\alpha}=\alpha\xi^{\alpha-1}(b-a).$
Consequently, we obtain the elegant representation
$\mathbb{E}\left[E(b)-E(a)\right]=\frac{\alpha\xi^{\alpha-1}}{\Gamma(1+\alpha)}(b-a),\quad\text{where }\xi\in(a,b).$
For any fixed $\varepsilon>0$, when we consider $a\geq\varepsilon$ (ensuring we avoid the singularity at the origin), we have $\xi^{\alpha-1}\leq\varepsilon^{\alpha-1}$ for all $\xi\geq\varepsilon$ (since $\alpha<1$). This yields the important upper bound
$\mathbb{E}\left[E(b)-E(a)\right]\leq C(b-a),$
where $C$ is a positive constant depending only on $\alpha$ and $\varepsilon$.
The moment bounds established in this section form the cornerstone of our subsequent analysis of numerical methods for time-changed SDEs. They provide rigorous control over the stochastic increments driven by the inverse subordinator $E(t)$ and directly influence the achievable convergence rates of our numerical scheme.
@@ 3 The strong convergence of EM method
4 Stability analysis of the EM method
Having established the strong convergence properties of the EM method, we now proceed to investigate its stability characteristics. We begin by giving the definitions of asymptotic mean square polynomial stability for the true solution and the numerical solution.
Definition 1.
The true solution $X(t)$ of the time-changed SDEs eq. 5 is said to be asymptotically polynomially stable in the mean square sense if there exists a positive constant $p$ such that
$\limsup_{t\to\infty}\frac{\log\mathbb{E}\left[|X(t)|^{2}\right]}{\log t}\leq-p.$
Definition 2.
The EM method defined by eq. 6 is said to be asymptotically polynomially stable in the mean square sense if there exists a positive constant $p$ such that
$\limsup_{n\to\infty}\frac{\log\mathbb{E}\left[|X_{n}|^{2}\right]}{\log t_{n}}\leq-p.$
To facilitate our stability analysis, we introduce the following assumptions on the coefficients of the time-changed SDEs.
Assumption 8.
There exists a positive constant $K$ such that for all $x\in\mathbb{R}^{d}$,
$|f(x)|\leq K|x|.$
Assumption 9.
There exist positive constants $L$ such that for all $x\in\mathbb{R}^{d}$,
$2\langle x,f(x)\rangle+|g(x)|^{2}\leq-L|x|^{2}.$
The condition in Assumption 9 is a strong dissipativity condition, which ensures that the combined effect of the drift term and the squared diffusion coefficient is negative definite. This guarantees that the system has a strong stabilizing tendency toward the origin, with the drift term’s contractive effect dominating any potential destabilizing influence from the diffusion term. This type of condition is commonly used in the stability analysis of SDEs and their numerical methods.
We now present our main theorem regarding the asymptotic stability properties of both the true solution and the EM method for time-changed SDEs.
Theorem 10.
Let $E$ be the inverse process of the subordinator $D$ with stability index $\alpha\in(0,1)$. Under Assumption 9, $f$ the initial condition $X_{0}$ is a constant, then the true solution of the time-changed SDEs eq. 5 is asymptotically polynomially stable in the mean square sense, i.e.,
$\limsup_{t\to\infty}\frac{\log\mathbb{E}\left[|X(t)|^{2}\right]}{\log t}\leq-\alpha.$
To prove this theorem, we require the following lemma from [8], which characterizes the asymptotic behavior of the inverse subordinator.
Lemma 11.
If the Bernstein function $\phi\in\textup{RV}_{\alpha}$ with $\alpha\in(0,1)$, then for any $\lambda>0$,
$\lim_{t\to\infty}\frac{\log\mathbb{E}_{D}\left[e^{-\lambda E(t)}\right]}{\log t}=-\alpha.$
Proof of Theorem 10.
The proof follows from similar arguments as in [21]. For completeness, we provide a detailed proof in B. ∎
Now, we proceed to analyze the stability properties of the EM method.
Theorem 12.
Let $E$ be the inverse process of the subordinator $D$ with stability index $\alpha\in(0,1)$. Under Assumptions 8 and 9, consider a family of uniform grids with step size $\Delta t\to 0$ and $t_{n}=n\Delta t\to\infty$. If the initial condition $X_{0}$ is a constant, then the EM method exhibits asymptotic polynomial stability in the mean-square sense with precise decay rate $\alpha$, i.e.,
$\limsup_{n\to\infty}\frac{\log\mathbb{E}\left[|X_{n}|^{2}\right]}{\log t_{n}}\leq-\alpha.$
Proof.
We begin by taking the square of the Euclidean norm on both sides of the iteration formula in eq. 6, which yields
$|X_{k+1}|^{2}$ $=|X_{k}|^{2}+|f(X_{k})|^{2}|\Delta E_{k}|^{2}+|g(X_{k})|^{2}|\Delta B_{k}|^{2}+2\langle X_{k},f(X_{k})\rangle\Delta E_{k}$
$\quad+2\langle X_{k},g(X_{k})\Delta B_{k}\rangle+2\langle f(X_{k})\Delta E_{k},g(X_{k})\Delta B_{k}\rangle.$
Since $(B(E(t)))_{t\geq 0}$ is a martingale, we have $\mathbb{E}_{B}[\Delta B_{k}]=0$. Taking $\mathbb{E}_{B}$ on both sides and noting that $\mathbb{E}_{B}[|\Delta B_{k}|^{2}]=\Delta E_{k}$, we obtain
$\mathbb{E}_{B}[|X_{k+1}|^{2}]$ $=\mathbb{E}_{B}[|X_{k}|^{2}]+\mathbb{E}_{B}[|f(X_{k})|^{2}]|\Delta E_{k}|^{2}+\mathbb{E}_{B}[|g(X_{k})|^{2}]\Delta E_{k}$
$\quad+2\mathbb{E}_{B}[\langle X_{k},f(X_{k})\rangle]\Delta E_{k}.$
Applying Assumptions 8 and 9, we arrive at the fundamental one-step inequality
$\mathbb{E}_{B}[|X_{k+1}|^{2}]$ $\leq(1-L\Delta E_{k}+K^{2}(\Delta E_{k})^{2})\mathbb{E}_{B}[|X_{k}|^{2}].$ (12)
The challenge in the subsequent analysis is that directly adding $(\Delta E_{k})^{2}$ into the exponent would yield a positive exponential term $\sum(\Delta E_{k})^{2}$, which could potentially dominate the polynomial decay induced by $L\sum\Delta E_{k}$. To circumvent this difficulty, we employ a localization argument based on stopping times.
For each $m\in\mathbb{N}$, define the stopping time
$\tau_{m}:=\inf\{t\geq 0:E(t)\geq m\}\quad(m\in\mathbb{N}),$
which increases to $\infty$ as $m\to\infty$ and satisfies $\tau_{m}<\infty$ almost surely. Since $E(\cdot)$ is continuous and bounded on the compact interval $0,\tau_{m}]$, its modulus of continuity
$\omega_{m}(\Delta t):=\sup_{0\leq s\leq\tau_{m},\,0\leq h\leq\Delta t}\big(E(s+h)-E(s)\big)$
satisfies $\omega_{m}(\Delta t)\to 0$ as $\Delta t\to 0$ by uniform continuity on compact sets. Consequently, for all $k$ such that the event $\{t_{k+1}\leq\tau_{m}\}$ holds, we have
$(\Delta E_{k})^{2}\leq\omega_{m}(\Delta t)\Delta E_{k}.$ (13)
Substituting (13) into (12) and using the elementary inequality $1+x\leq e^{x}$, we obtain
$\mathbb{E}_{B}\left[|X_{k+1}|^{2}\mathbf{1}_{\{t_{k+1}\leq\tau_{m}\}}\right]$ $\leq\exp\big(-(L-K^{2}\omega_{m}(\Delta t))\Delta E_{k}\big)\mathbb{E}_{B}\left[|X_{k}|^{2}\mathbf{1}_{\{t_{k}\leq\tau_{m}\}}\right].$
Iterating this inequality from $k=0,\ldots,n-1$ yields
$\mathbb{E}_{B}\left[|X_{n}|^{2}\mathbf{1}_{\{t_{n}\leq\tau_{m}\}}\right]\leq|X_{0}|^{2}\exp\big(-(L-K^{2}\omega_{m}(\Delta t))E(t_{n}\wedge\tau_{m})\big).$ (14)
On the event $\{t_{n}\leq\tau_{m}\}$, we have $t_{n}\wedge\tau_{m}=t_{n}$. Taking $\mathbb{E}_{D}$ on both sides of (14) and using $\mathbf{1}_{\{t_{n}\leq\tau_{m}\}}\leq 1$, we obtain
$\mathbb{E}\!\left[|X_{n}|^{2}\mathbf{1}_{\{t_{n}\leq\tau_{m}\}}\right]\leq|X_{0}|^{2}\,\mathbb{E}_{D}\!\left[\exp\!\big(-(L-K^{2}\omega_{m}(\Delta t))\,E(t_{n})\big)\right].$
Since $\omega_{m}(\Delta t)\to 0$ as $\Delta t\to 0$ for each fixed $m$, there exists $N(m)\in\mathbb{N}$ such that for all $n\geq N(m)$,
$K^{2}\omega_{m}(\Delta t)\leq\frac{L}{2}.$ (15)
Then, by applying (15), we obtain for all $n\geq N(m)$,
$\mathbb{E}\left[|X_{n}|^{2}\mathbf{1}_{\{t_{n}\leq\tau_{m}\}}\right]$ $\leq\mathbb{E}\left[|X_{0}|^{2}\right]\mathbb{E}_{D}\left[e^{-\frac{L}{2}E(t_{n})}\right].$ (16)
By Lemma 11, taking $\lambda=\frac{L}{2}$ and letting $n\to\infty$, we deduce that for any fixed $m$,
$\limsup_{n\to\infty}\frac{\log\mathbb{E}\left[|X_{n}|^{2}\mathbf{1}_{\{t_{n}\leq\tau_{m}\}}\right]}{\log t_{n}}$ $\leq\limsup_{n\to\infty}\frac{\log\mathbb{E}_{D}\left[e^{-\frac{L}{2}E(t_{n})}\right]}{\log t_{n}}=-\alpha.$
To remove the localization, we invoke the monotone convergence theorem. Since $\tau_{m}\to\infty$ almost surely and $\mathbf{1}_{\{t_{n}\leq\tau_{m}\}}\to 1$ as $m\to\infty$, we have
$\mathbb{E}[|X_{n}|^{2}]=\lim_{m\to\infty}\mathbb{E}\left[|X_{n}|^{2}\mathbf{1}_{\{t_{n}\leq\tau_{m}\}}\right].$
Consequently,
$\limsup_{n\to\infty}\frac{\log\mathbb{E}[|X_{n}|^{2}]}{\log t_{n}}\leq-\alpha,$
which completes the proof. ∎
@@ 5 Numerical simulations
For the stability analysis, we simulate $10^{4}$ sample paths up to the terminal time $T=40$ with a step size of $\Delta t=2^{-13}$ and stability index $\alpha=0.6,0.7,0.8$. This fine discretization allows us to accurately capture the long-term behavior of the example while ensuring numerical stability according to the conditions established in Theorems 10 and 12.

### v2

# Parameter-related strong convergence rate of an Euler’s type method for time-changed stochastic differential equations

@@ Abstract
An Euler’s type method with the equidistant step size is proposed for a class of time-changed stochastic differential equations driven by the multiplicative noise and the strong convergence rate that is related to the parameter of the time changing process is obtained. Such a observation of the convergence rate is significantly different from those existing results that employ methods with the random step size. Numerical simulations are provided to demonstrate the theoretical results.
@@
time-changed stochastic differential equations, inverse subordinator, equidistant-step Euler-Maruyama method, strong convergence
@@ 1 Introduction
where $\mathcal{L}_{\text{FP}}=-\frac{\partial}{\partial x}\left(f(x)\cdot\right)+\frac{1}{2}\frac{\partial^{2}}{\partial x^{2}}\left(g(x)^{2}\cdot\right)$. In contrast, when considering subdiffusion phenomena via time-changed SDEs, as rigorously established by [5, 4], the probability density function of a time-changed process $X(E(t))$ (to be formally defined subsequently) is governed by a fractional Fokker-Planck equation(FFPE)
@@
Given the analytical complexity of time-changed SDEs, numerical methods play a crucial role in their practical implementation. The first systematic investigation of numerical methods for time-changed SDEs was conducted by [6], who considered the following time-changed SDEs
@@
Following this pioneering work, [7] investigated a class of time-changed SDEs where the inverse subordinator $E(t)$ in the drift and diffusion coefficients of eq. 1 is replaced with ordinary time $t$. In this modified framework, the duality principle no longer applies, necessitating the development of alternative analytical techniques. Using a Gronwall-like inequality specific to time-changed integrals, they demonstrated that the EM method maintains a strong convergence rate of $1/2$ in this setting as well.
@@
In this paper, we consider the following time-changed SDEs
@@
where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times m}$ are coefficient functions satisfying global Lipschitz and linear growth conditions, $E(t)$ denotes the inverse subordinator, and $B(E(t))$ represents the $m$-dimensional time-changed Brownian motion.
Our contributions are twofold
@@
The remainder of this paper is organized as follows: Section 2 introduces the mathematical preliminaries, including key properties of subordinators and their inverses, and establishes important lemmas used in our subsequent analysis. Section 3 presents the fundamental assumptions, develops the equidistant-step EM scheme, and proves our main convergence theorem. Section 4 presents numerical experiments that validate our theoretical findings and illustrate the effectiveness of the proposed method.
@@ 2 Preliminaries
To numerically simulate the inverse subordinator $E$, we employ the method developed by [14, 15]. Given a time horizon $T>0$ and an equidistant step size $\delta>0$, we first generate a sample path of $D$ over $[0,T]$ by setting $D(0)=0$ and recursively computing
@@
The equidistant-step EM method presented in eq. 4 has been previously utilized by [16] in their investigation of anomalous diffusion processes, which provide stochastic representations for multidimensional FFPE. Their approach employed sophisticated techniques from Malliavin calculus to derive unbiased density formulas for the probability density functions of these anomalous diffusion processes. This methodology yielded remarkably accurate numerical approximations that successfully captured essential characteristics of the underlying processes, including the distinctive cusp formations observed in density profiles. Despite these promising applications, a rigorous mathematical analysis concerning the convergence properties of this numerical scheme in the context of time-changed SDEs has not been thoroughly established in the literature. The present work endeavors to fill this significant gap in the theoretical understanding of numerical methods.
@@
The following lemma from [17] provides precise bounds on the expected values of multiple integrals involving the inverse subordinator. This result constitutes the cornerstone of our convergence rate analysis, establishing the direct relationship between the time step size and the expectation of multiple integrals driven by increments of $E(t)$.
@@
Let $E=\{E(t),t\geq 0\}$ be the inverse $\alpha$-stable subordinator with index $0<\alpha<1$. Then for all $0<a<b<\infty$ and all integers $n\geq 1$, the following bounds hold
$\left(\frac{b-a}{b}\right)^{1-\alpha}\frac{n!(b-a)^{n\alpha}}{\Gamma((n-1)\alpha+2)\Gamma(\alpha)}\leq\mathbb{E}\left[|E(b)-E(a)|^{n}\right]\leq\left(\frac{b-a}{b}\right)^{1-\alpha}\frac{n!(b-a)^{n\alpha}}{\Gamma(n\alpha+1)}.$
Lemma 3 provides the precise asymptotic estimates that are fundamental to our subsequent convergence rate analysis. The upper bound in this lemma reveals that, for fixed $b$ and sufficiently small $\Delta t$ with $b-a=\Delta t$, the $n$-th moment of the increment E(b)E(a)E(b)-E(a) satisfies
$\mathbb{E}\left[|E(b)-E(a)|^{n}\right]\leq C(\Delta t)^{1+(n-1)\alpha},$
where the constant $C$ depends on $b$ but remains independent of the step size $\Delta t$. This crucial estimate demonstrates how the parameter $\alpha$ governs the growth rate of moments for increments of the inverse subordinator, and plays a pivotal role in establishing the convergence order of the numerical method in the following section.